n belongs to natural number
find a1,a2,a3,....,an belonging to natural number such that a1a2,a2a3,a3a4,........,an-1an are in AP
 

Grade:12th pass

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

a1 = a+d

a2= a+2d

a3= a+3d

and so on

2(a+d)(a+2d) = (a*(a+d))+(a+2d)(a+3d)

also the next ie 2(3rd term ) = (2nd term + 4thterm)

using 2 equations a and d can be identified

and the question can be solved

Regards